OpenStudy (mayankdevnani):

A metal cube of edge 12 cm is melted and formed into three smaller cubes. If the edges of two smaller cubes are 6 cm and 8 cm . Find the edge of the third smaller cube.(Assume that there is no loss of metal during melting).

OpenStudy (shamim):

12*12*12=6*6*6+8*8*8+l*l*l

OpenStudy (mayankdevnani):

what???plz explain me

OpenStudy (shamim):

u hv to solve this equation

OpenStudy (shamim):

ok

OpenStudy (shamim):

the volume of larger cube is equal to the total volume of those small three cube

OpenStudy (anonymous):

\[V_{initial}=V_{final}\] \

OpenStudy (shamim):

ya u r righr

OpenStudy (mayankdevnani):

ok

OpenStudy (shamim):

do u wnt to get the solution of my equation

OpenStudy (mayankdevnani):

12*12*12=6*6*6+8*8*8+l*l*l